Word problems are not my stong suit...please help.
A plane flies 720 mi against a steady 30-mi/hr headwind and then returns to the same point with the wind. If the entire trip takes 10 h., what is the plane's speed in still air?

Answer by checkley75(3666)

720/(x-30)+720/(x+30)=10
[720(x+30)+720(x-30)]/(x-30)(x+30=10
720x+600+720x-600/(x^2-900)=10
1440x/(x^2-900)=10 cross multiply
1440x=10x^2-9000
10x^2-1440x-9000=0
10(x^2-144-900)=0
10(x-150)(x+6)=0
x-150=0
x=150 mph for the plane.
proof
720/(150-30)+720/(150+30)=10
720/120+720/180=10
6+4=10
10=10